OpenStudy (anonymous):

factor completely? a^4-50a^2b^2+49b^4

jimthompson5910 (jim_thompson5910):

Let \(\Large m = a^2\) and \(\Large n = b^2\) \[\Large a^4-50a^2b^2+49b^4\] \[\Large (a^2)^2-50a^2b^2+49(b^2)^2\] \[\Large m^2-50mn+49n^2\] \[\Large (m-49n)(m-n)\] \[\Large (a^2-49b^2)(a^2-b^2)\] \[\Large (a^2-(7b)^2)(a^2-b^2)\] \[\Large (a-7b)(a+7b)(a^2-b^2)\] \[\Large (a-7b)(a+7b)(a-b)(a+b)\] ------------------------------------------------------- So \[\Large a^4-50a^2b^2+49b^4\] completely factors to \[\Large (a-7b)(a+7b)(a-b)(a+b)\]